Two blocks A(3 kg ) and B(6kg) are connected by a spring stiffness 512N/m and placed on smooth horizontal surface. Initially the spring has equilibrium length. Velocity 1.8m/s and 2.2m/s are imparted to A and B in opposite direction. The maximum extension of the spring will be______cm

Grade:11

1 Answers

Karan Gupta
6 years ago
At maximum extension of the spring, velocities of both blocks will be same.
 
Conservation of momentum:
6*2.2 – 3*1.8 = (3+6)*V
V = 0.867 m/s
 
Conservation of energy:
1/2*3*1.8*1.8 + 1/2*6*2.2*2.2 = 1/2*(6+3)*0.867*0.867 + 1/2*512* X*X
X = 0.25 m = 25 cm